OpenStudy (shamil98):

find the exact length of the curve. x = tsint , y = tcost i've come up with \[\int\limits_{0}^{1} \sqrt{2t^2+1}\] after doing dx/dt and dy/dt and squaring under the square root symbols and using trig identities how do i proceed to do this without a calculator now? or have i messed up along the way? \[dx/dt = sint + tcost \] \[dy/dt = cost-tsint\] \[L = \int\limits_{0}^{1}\sqrt{(sint + tcost)^2 +(cost-tsint)^2 }\]

OpenStudy (shamil98):

i think i've found my mistake one sec

OpenStudy (shamil98):

came up with this i i thought it was t^2(cos^2 + sin^2) by its this instead \[L = \int\limits_{0}^{1} \sqrt{t^2(\cos^2t*\sin^2 t) + 1}\]